if yhat=-114 +2.88 x-sub2 +1.86 x-sub-2
how do I know what x is equal to ?
x-sub-1 is the female athlete's height (in inches) and x-sub-2 is the female athlete's percent body fat, measured as x2%.
what is showing on th multiple regression equation to predict the total body weight for a female athlete who is 67 inches tall and has 28% body fat.
I cannot figure out what the value for X is equal to.
I do know that I will have to subtract two numbers lower than x from 114 and subtracting 2 from x i will add sum to 1.86
Does tis sound right
I am 63 years old and this is my last class before a long awaited degree :(



Answer by Boreal(15235)

y hat=2.88 x1+1.86 x2-114; x1=female's height in inches (note editing, is this what you meant?) and x2= body fat in %
2.88(67)+1.86(28)-114=
192.96+52.08-114=245.04-114=131.04 pounds.
y hat is the predicted body weight as a function of two variables (height and body fat) minus a constant 114.
I think this is what you meant, and the answer would appear to make sense, but re-submit if this isn't what you wanted.